OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. a.Set up the system of equations that needs to be solved to determine how many of each type of phone were sold. Give a clear definition of the variables in the system. b.Solve the system of equations, showing clearly how the solution was determined, and state the results clearly in light of the real-world situation. c.Verify your results of the 2 linear e

hero (hero):

Let's go back to vyew

hero (hero):

I don't know why no one else helps you with these

OpenStudy (anonymous):

lol their long questions i think thats why rofl

hero (hero):

Well, I have a surprise for you

OpenStudy (anonymous):

yeah?

hero (hero):

Come to vyew and I'll tellyou about it

OpenStudy (anonymous):

lol ok c.Verify your results of the 2 linear equations by graphing in the desired graphing program and paste the graph in your assignment document (edit/copy image). d.Explain how the results are verified by the graph.

hero (hero):

x = Cell Phone Model A4 y = Smart Phone Z20 a. Set up the system: x + y = 51 35x + 50y = 2220 b. Solve the system: y = 51 - x y = (2220 - 35x)/50 y = y 50(51-x) = 2220 - 35x 2550 - 50x = 2220 - 35x 30 = 15x 30/15 = x 2 = x y = 49 c. Graph these two equations: y = 51 - x y = (2220 - 35x)/50 The two lines intersect at (x,y) = (2,49) d. The results are verified by the graph because the same solution obtained by algebraic methods are also found using the graphing methods

OpenStudy (anonymous):

very Nice!!! Thanks Hero!

hero (hero):

You assume I still have your email. I don't

OpenStudy (anonymous):

lol doh! i appologize about that. my e-mail is romieislam@live.com